# A star is born: Explosive Crump-Mode-Jagers branching processes in random environment

### v1

@@ 2.2.1. Super-linear preferential attachment with fitness
We proceed by studying a family of preferential attachment trees (with fitness) for which we can verify the conditions provided in Corollary 2.7. Most notably, we significantly enlarge the class of models that fit within the framework of super-linear preferential attachment trees with fitness for which Corollary 2.7 holds. Indeed, we extend this from only a handful of examples to a wide class of models. These consist of a family of multiplicative fitness functions. That is, vertex-weights take values in $S=\mathbb{R}_{+}$, and $f(i,w)=g(w)s(i)$ for some functions $g\colon\mathbb{R}_{+}\to(0,\infty)$ and $s\colon\mathbb{N}_{0}\to(0,\infty)$. To satisfy Assumption ($w^{*}$), we assume that
@@
Our main interest is functions $s$ that grow barely faster than linear, for which the summability condition in (2.11) is only just sastisfied. For example, $s(n)=n\log(n+2)\log\log(n+3)^{\sigma}$, with $\sigma>1$. Earlier work by the author and Iyer in [10] is only able to deal with cases for which $s(n)$ grows fast enough, that is, faster than $n(\log n)^{\alpha}$ with $\alpha>2$. Here, we extend this to a much wider range of functions that grow slower.
@@
Suppose that $f(i,w)=g(w)s(i)$, where $s$ satisfies Assumption 2.10 and $g$ and $s$ satisfy (2.11). Suppose there exists a sequence $(k_{n})_{n\in\mathbb{N}}$ and constants $\varepsilon>0,n_{0}\in\mathbb{N}$, such that
$\mathbb{P}\!\left(g(W)>k_{n}\right)\leq n^{-(1+\varepsilon)},\qquad\text{for all }n\geq n_{0}.$ (2.15)
Then, if there exists $\delta\in(0,\varepsilon)$ such that
$\lim_{n\to\infty}\frac{\mu_{(\delta\log(n)/(\mu_{n}k_{n}))^{1/\beta}}}{\mu_{n}k_{n}}=\infty,$ (2.16)
the limiting infinite tree $T_{\infty}$ contains a unique infinite star and no infinite path, almost surely.
@@ 2.3. Examples
[…]
@@
Consider the four examples $(i)$-$(iv)$. The infinite tree $T_{\infty}$ almost surely contains a unique infinite star and no infinite path when
@@
When the above inequalities are reversed, almost surely $T_{\infty}$ contains a unique infinite path and no infinite star.
@@ 2.3.1. Super-linear preferential attachment without fitness
$\lim_{n\to\infty}\frac{\mu_{(\log(n)\mu_{n}^{-1})^{1/\beta}}}{\mu_{n}}=\infty.$ (2.20)
Then, the limiting infinite tree $T_{\infty}$ contains a unique vertex with infinite degree and no infinite path, almost surely.
@@ 3. Proofs of main results
$\sum_{\begin{subarray}{c}a:|a|=m\\ a\text{ is $a_{1}$-conservative}\end{subarray}}\!\!\!\!\!\!\!\!\!\!\!\!\!\mathbb{P}\!\left(a\text{ has at least $a_{1}$ children before }\varnothing\text{ explodes}\right)\leq\zeta^{m-1}\mathcal{M}_{\lambda_{a_{1}}}(Y_{a_{1}})\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{a_{1}};W)\right].$ (3.1)
@@
$\mathbb{P}\!\left(a\text{ has at least $a_{1}$ children before }\varnothing\text{ explodes}\right)=\mathbb{P}\bigg{(}\!\mathcal{P}_{a_{1}}(a)+\!\sum_{j=2}^{m}\mathcal{P}_{a_{j}}(a_{|_{j-1}})\leq\!\!\!\!\sum_{k=a_{1}+1}^{\infty}\!\!\!\!X(k)\!\bigg{)}.$ (3.2)
@@
$\mathbb{P}\bigg{(}\!\mathcal{P}_{a_{1}}(a)+\!\sum_{j=2}^{m}\mathcal{P}_{a_{j}}(a_{|_{j-1}})\leq Y_{a_{1}}\!\bigg{)}\leq\mathcal{M}(\lambda_{a_{1}}(Y_{a_{1}}))\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{a_{1}};W)\right]\prod_{j=2}^{m}\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{a_{j}};W)\right],$ (3.3)
@@
$\sum_{\begin{subarray}{c}a:|a|=m\\ a\text{ $a_{1}$-conservative}\end{subarray}}{}$ $\mathbb{P}\!\left(a\text{ has at least $a_{1}$ children before }\varnothing\text{ explodes}\right)$ (3.4)
@@
[…]
@@
$\sum_{n=1}^{a_{1}}\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{n};W)\right]<\zeta.$
@@
$\sum_{n=1}^{\infty}\mathbb{E}\left[\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n};W)\right]<\infty$ (3.5)
@@
$\sum_{n=N}^{a_{1}}\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{n};W)\right]<\sum_{n=N}^{\infty}\mathbb{E}\left[\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n};W)\right]<\frac{\eta}{2},$ (3.6)
@@
$\lim_{a_{1}\to\infty}\sum_{n=1}^{N-1}\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{n};W)\right]=\sum_{n=1}^{N-1}\mathbb{P}\!\left(\mathcal{P}_{n}=0\right)=\sum_{n=1}^{N-1}\mathbb{P}\!\left(X(1)=\cdots=X(n)=0\right).$ (3.7)
@@
$\mathbb{E}\left[\sup\{k:X(k)=0\}\right]=\sum_{n=1}^{\infty}n\mathbb{P}\!\left(X(1)=\cdots=X(n)=0\right)<1-\xi.$ (3.8)
Hence, the right-hand side of (3.7) is at most $1-\xi$ for any $N\in\mathbb{N}$ as well. So, we first take $\eta<\xi$ and $N$ large enough so that we have the upper bound in (3.6). Then, we take $K\geq N$ sufficiently large, so that for all $a_{1}\geq K$,
$\sum_{n=1}^{N-1}\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{\ell};W)\right]<1-\xi/2.$ (3.9)
@@
$\sum_{n=1}^{a_{1}}\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{\ell};W)\right]<1-\xi/2=:\zeta.$ (3.10)
@@
The above lemma provides an upper bound for the probability of the event that a vertex $a$ explodes before the root of $\mathcal{U}_{\infty}$, in the case that $a$ is $a_{1}$-conservative. When $a$ does not satisfy this condition, we can view $a$ as a concatenation of a number of conservative sequences. That is, we write $a=\overline{b}_{1}\cdots\overline{b}_{\ell}$, where $\overline{b}_{i}=b_{i,1}\ldots b_{i,m_{i}}$ for each $i\in[\ell]$ and for some $\ell\in\mathbb{N},(m_{i})_{i\in[\ell]}\in\mathbb{N}^{\ell}$, and $(b_{i,j})_{i\in[\ell],j\in[m_{i}]}$, such that $\overline{b}_{i}$ is $b_{i,1}$-conservative for each $i\in[\ell]$. By the independence of birth processes of distinct individuals (or in fact, the independence of disjoint sub-trees), we are able to apply Lemma 3.1 to each conservative sequence in the concatenation to arrive at a bound for the expected number of individuals that explode before all its ancestors.
@@
As explained before the proposition statement, we think of sequences $a\in\mathcal{U}_{\infty}$ as a concatenation of conservative sequences. Let $a=a_{1}\ldots a_{m}$ be a sequence of length $m\in\mathbb{N}$, and assume that there exist $k\in[m]$ and indices $I_{1}<I_{2}<\ldots<I_{k}$ such that $a_{1}=:a_{I_{1}}<a_{I_{2}}<\ldots<a_{I_{k}}$. That is, the $I_{j}$ are the indices of the running maxima of $a$. We also set $I_{k+1}:=m+1$ and $a_{0}=\varnothing$. We think of $a$ as a concatenation of the conservative sequences $a_{I_{j}}\cdots a_{I_{j+1}-1}$, with $j\in[k]$. These sub-sequences can be seen as corresponding to an $a_{I_{j}}$-conservative individual, rooted at $a_{I_{j}-1}$, for each $j\in[k]$. We can thus apply Lemma 3.1 to these sub-sequences.
@@
[…]
@@
(2.1)${{\leq}}\prod_{\ell=1}^{k}\mathbb{P}\bigg{(}\mathcal{P}_{a_{I_{\ell}}}(a_{|_{I_{\ell+1}-1}})+\sum_{j=I_{\ell}}^{I_{\ell+1}-2}\mathcal{P}_{a_{j+1}}(a_{|_{j}})\leq Y^{(a_{1}\cdots a_{I_{\ell}-1})}_{a_{I_{\ell}}}\,\bigg{|}\,W_{\varnothing},W_{a_{1}},W_{a_{1}a_{2}},\ldots,W_{a}\bigg{)}$ )
@@
We now let $d_{j}:=I_{j+1}-I_{j}-1$ for $j\in[k-1]$ and $d_{k}:=m-I_{k}$ denote the number of entries between the running maxima of $a$. We can then define, for $(d_{j})_{j\in[k]}\in\mathbb{N}_{0}^{k}$ (with $[0]:=\varnothing$),
$\mathscr{P}_{k}(a_{I_{1}},a_{I_{2}},\ldots,a_{I_{k}},d_{1},\ldots,d_{k}):=\{a\in\mathcal{U}_{\infty}:\text{ For all }j\in\{1,\ldots,k\}\text{ and all }i\in[d_{j}],\ a_{I_{j}+i}\in[a_{I_{j}}]\}$ (3.15)
@@
$\prod_{\ell=1}^{k}\!\!\!\!\sum_{\begin{subarray}{c}a^{(\ell)}:|a^{(\ell)}|=d_{\ell}\\ \text{$a^{(\ell)}$ $a_{I_{\ell}}$-conservative}\end{subarray}}\!\!\!\!\!\!\!\!\!\!\!\mathbb{P}\!\left(\widetilde{E}_{a,\ell}\right)\leq\prod_{\ell=1}^{k}\zeta^{d_{\ell}}\mathcal{M}_{\lambda_{a_{I_{\ell}}}}(Y_{a_{I_{\ell}}})\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{I_{\ell}}}}(\mathcal{P}_{a_{I_{\ell}}};W)\right].$ (3.18)
We can take out the factors $\zeta^{d_{\ell}}$ and use that the $d_{\ell}$ sum to $m-k$. Using this in (3.16) yields
$\sum_{m=1}^{\infty}{}$ $\sum_{k=1}^{m}\sum_{a_{I_{k}}>\ldots>a_{I_{1}}>K^{\prime}}\zeta^{m-k}\sum_{\begin{subarray}{c}(d_{j})_{j\in[k]}\in\mathbb{N}_{0}^{k}\\ \sum_{\ell=1}^{k}d_{\ell}=m-k\end{subarray}}\prod_{\ell=1}^{k}\mathcal{M}_{\lambda_{a_{I_{\ell}}}}(Y_{a_{I_{\ell}}})\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{I_{\ell}}}}(\mathcal{P}_{a_{I_{\ell}}};W)\right]$ (3.19)
$\leq\sum_{m=1}^{\infty}\sum_{k=1}^{m}\binom{m-1}{k-1}\zeta^{m-k}\sum_{a_{I_{1}}>K^{\prime}}\cdots\sum_{a_{I_{k}}>K^{\prime}}\prod_{\ell=1}^{k}\mathcal{M}_{\lambda_{a_{I_{\ell}}}}(Y_{a_{I_{\ell}}})\mathbb{E}\left[\mathcal{L}_{\lambda_{a_{I_{\ell}}}}(\mathcal{P}_{a_{I_{\ell}}};W)\right]$
$=\sum_{m=1}^{\infty}\sum_{k=1}^{m}\binom{m-1}{k-1}\zeta^{m-k}\left(\sum_{n>K^{\prime}}\mathcal{M}_{\lambda_{n}}(Y_{n})\mathbb{E}\left[\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n};W)\right]\right)^{k}.$
@@
$\mathcal{U}_{L}:=\left\{u\in\mathcal{U}_{\infty}:u_{i}\leq L,i\in[|u|]\right\}\cup\{\varnothing\}.$
@@
Let $(\mathscr{T}_{t})_{t\geq 0}$ be a $(X,W)$-CMJ branching process satisfying Condition (c) of Assumption 2.1. Then, almost surely, for all $t\in(0,\infty)$, we have $|\mathscr{T}_{t}\cap\mathcal{U}_{L}|<\infty$.
Recall that, for an $(X,W)$-Crump-Mode-Jagers branching process $(\mathscr{T}_{t})_{t\geq 0}$, we have
@@
Now, since $\tau_{\infty}<\infty$ almost surely, we infer from Proposition 3.3 with $L=K^{\prime}$, that |$|\{u\in\mathcal{U}_{K^{\prime}}:\mathcal{B}(u)\leq\tau_{\infty}\}|<\infty$ almost surely. Therefore, by (3.21), the set
@@
The desired result follows by verifying the conditions on $\lambda_{n}$ and the vertex-weights in Corollary 2.7. We start by constructing a sequence $(\lambda_{n})_{n\in\mathbb{N}}$ such that there exists $N\in\mathbb{N}$ for which we have $\lambda_{n}<f(i,w)$ for all $i\geq n\geq N$. We recall that there exist $\varepsilon>0$, $n_{0}\in\mathbb{N}$, and a sequence $(k_{n})_{n\in\mathbb{N}}$ such that $\mathbb{P}\!\left(g(W)>k_{n}\right)\leq n^{-(1+\varepsilon)}$ for all $n\geq n_{0}$. Then, we fix $\delta\in(0,\varepsilon)$, recall $\mu_{n}$ from (2.7), and set
@@
[…]

### v2

@@ 2.2.1. Super-linear preferential attachment with fitness
We proceed by studying a family of preferential attachment trees (with fitness) for which we can verify the conditions provided in Corollary 2.7. In doing so, we extend the number of models for which it is known that $T_{\infty}$ contains a unique node of infinite degree from a handful of examples to larger, but more particular, class. This class consists of a family of multiplicative fitness functions. That is, vertex-weights take values in $S=\mathbb{R}_{+}$, and $f(i,w)=g(w)s(i)$ for some functions $g\colon\mathbb{R}_{+}\to(0,\infty)$ and $s\colon\mathbb{N}_{0}\to(0,\infty)$. To satisfy Assumption ($w^{*}$), we assume that
@@
Our main interest is functions $s$ that grow barely faster than linear, for which the summability condition in (2.11) is only just satisfied. For example, $s(n)=n\log(n+2)\log\log(n+3)^{\sigma}$, with $\sigma>1$. Earlier work by the author and Iyer in [10] is only able to deal with cases for which $s(n)$ grows fast enough, that is, faster than $n(\log n)^{\alpha}$ with $\alpha>2$. Here, we extend this to a much wider range of functions that grow slower.
@@
Suppose that $f(i,w)=g(w)s(i)$, where $s$ satisfies Assumption 2.10 and $g$ and $s$ satisfy (2.11). Suppose there exists a sequence $(k_{n})_{n\in\mathbb{N}}$ and constants $\varepsilon>0,\delta\in(0,\varepsilon)$, and $n_{0}\in\mathbb{N}$, such that both
$\mathbb{P}\!\left(g(W)>k_{n}\right)\leq n^{-(1+\varepsilon)},\qquad\text{for all }n\geq n_{0},$ (2.15)
and
$\lim_{n\to\infty}\frac{\mu_{(\delta\log(n)/(\mu_{n}k_{n}))^{1/\beta}}}{\mu_{n}k_{n}}=\infty$ (2.16)
are satisfied, where $\beta$ is as in Assumption 2.10. Then, the limiting infinite tree $T_{\infty}$ contains a unique infinite star and no infinite path, almost surely.
@@ 2.3. Examples
$({}$ $i)$ $s(i)=(i+1)(\log(i+2))^{\sigma},\quad$ $\mathbb{P}\!\left(W\geq x\right)=\Theta\big(\mathrm{e}^{-x^{\kappa}}\big),$ (2.17)
$({}$ $ii)$ $s(i)=(i+1)\log(i+2)\exp((\log\log(i+3))^{\nu}),\quad$ $\mathbb{P}\!\left(W\geq x\right)=\Theta(\exp\big(-\mathrm{e}^{(\log x)^{\gamma}}\big)\big),$
$({}$ $iii)$ $s(i)=(i+1)\log(i+2)(\log\log(i+3))^{\sigma},\quad$ $\mathbb{P}\!\left(W\geq x\right)=\Theta\big(\exp\big(-\mathrm{e}^{x^{\kappa}}\big)\big),$
@@
Consider the four examples (i)(i)-$(iv)$. The infinite tree $T_{\infty}$ almost surely contains a unique infinite star and no infinite path when
@@
When the above inequalities are reversed ((i.e. when changing the >> to a <)<), almost surely $T_{\infty}$ contains a unique infinite path and no infinite star.
@@ 2.3.1. Super-linear preferential attachment without fitness
$\lim_{n\to\infty}\frac{\mu_{(\log(n)\mu_{n}^{-1})^{1/\beta}}}{\mu_{n}}=\infty,$ (2.20)
where $\beta$ is as in Assumption 2.10. Then, the limiting infinite tree $T_{\infty}$ contains a unique vertex with infinite degree and no infinite path, almost surely.
@@ 3. Proofs of main results
[…]
@@
[…]
@@
$\mathbb{P}\bigg(\!\mathcal{P}_{a_{1}}(a)+\!\sum_{j=2}^{m}\mathcal{P}_{a_{j}}(a_{|_{j-1}})\leq Y_{a_{1}}\!\bigg)\leq\mathcal{M}(\lambda_{a_{1}}(Y_{a_{1}}))\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{a_{1}})\prod_{j=2}^{m}\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{a_{j}}),$ (3.3)
@@
$\leq{}$ $\sum_{a_{2}=1}^{a_{1}}\sum_{a_{3}=1}^{a_{1}}\cdots\sum_{a_{m}=1}^{a_{1}}\mathcal{M}(\lambda_{a_{1}}(Y_{a_{1}}))\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{a_{1}})\prod_{j=2}^{m}\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{a_{j}})$
$={}$ $\mathcal{M}(\lambda_{a_{1}}(Y_{a_{1}}))\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{a_{1}})\left(\sum_{n=1}^{a_{1}}\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{n})\right)^{m-1}.$
@@
$\sum_{n=1}^{a_{1}}\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{n})<\zeta.$
@@
$\sum_{n=1}^{\infty}\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n})<\infty$ (3.5)
@@
$\sum_{n=N}^{a_{1}}\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{n})<\sum_{n=N}^{\infty}\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n})<\frac{\eta}{2},$ (3.6)
@@
$\lim_{a_{1}\to\infty}\sum_{n=1}^{N-1}\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{n})=\sum_{n=1}^{N-1}\mathbb{P}\!\left(\mathcal{P}_{n}=0\right)=\sum_{n=1}^{N-1}\mathbb{P}\!\left(X(1)=\cdots=X(n)=0\right).$ (3.7)
@@
$\mathbb{E}\left[\sup\{k:X(1)=\cdots=X(k)=0\}\right]=\sum_{n=1}^{\infty}\mathbb{P}\!\left(X(1)=\cdots=X(n)=0\right)<1-\xi.$ (3.8)
Hence, the right-hand side of (3.7) is at most $1-\xi$ for any $N\in\mathbb{N}$. So, we first take $\eta<\xi$ and $N$ large enough so that we have the upper bound in (3.6). Then, we take $K\geq N$ sufficiently large, so that for all $a_{1}\geq K$,
$\sum_{n=1}^{N-1}\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{\ell})<1-\xi/2.$ (3.9)
@@
$\sum_{n=1}^{a_{1}}\mathcal{L}_{\lambda_{a_{1}}}(\mathcal{P}_{\ell})<1-\xi/2=:\zeta.$ (3.10)
@@
The above lemma provides an upper bound for the probability of the event that a vertex $a$ explodes before the root of $\mathcal{U}_{\infty}$, in the case that $a$ is $a_{1}$-conservative. When $a$ does not satisfy this condition, we can view $a$ as a concatenation of a number of conservative sequences. That is, we write $a=\overline{b}_{1}\cdots\overline{b}_{\ell}$, where $\overline{b}_{i}=b_{i,1}\ldots b_{i,m_{i}}$ for each $i\in[\ell]$ and for some $\ell\in\mathbb{N},(m_{i})_{i\in[\ell]}\in\mathbb{N}^{\ell}$, and $(b_{i,j})_{i\in[\ell],j\in[m_{i}]}$, such that $\overline{b}_{i}$ is $b_{i,1}$-conservative for each $i\in[\ell]$. By the independence of birth processes of distinct individuals (or in fact, the independence of disjoint sub-trees), we are able to apply Lemma 3.1 to each conservative sequence in the concatenation to arrive at a bound for the expected number of individuals that explode before all their ancestors.
@@
As explained before the proposition statement, we think of sequences $a\in\mathcal{U}_{\infty}$ as a concatenation of conservative sequences. Let $a=a_{1}\ldots a_{m}$ be a sequence of length $m\in\mathbb{N}$, and assume that there exist $k\in[m]$ and indices $1:=I_{1}<I_{2}<\ldots<I_{k}$ such that $a_{1}=:a_{I_{1}}<a_{I_{2}}<\ldots<a_{I_{k}}$ and $a_{i}\leq a_{I_{j}}$ for all $i\in\{I_{j}+1,\ldots,I_{j+1}-1\}$ and $j\in[k]$ (with $I_{k+1}:=m+1$). That is, the $I_{j}$ are the indices of the running maxima of $a$. We also set $I_{k+1}:=m+1$ and $a_{0}=\varnothing$. We think of $a$ as a concatenation of the conservative sequences $a_{I_{j}}\cdots a_{I_{j+1}-1}$, with $j\in[k]$. These sub-sequences can be seen as corresponding to an $a_{I_{j}}$-conservative individual, rooted at $a_{I_{j}-1}$, for each $j\in[k]$. We can thus apply Lemma 3.1 to these sub-sequences.
@@
$E_{a}$={} $\{a\text{ explodes}\text{ before any of its ancestors explodes}\}$ (3.12)
@@
$={}$ $\bigcap_{\ell=1}^{k}\bigg\{\mathcal{P}_{a_{I_{\ell}}}(a_{|_{I_{\ell+1}-1}})+\sum_{j=I_{\ell}}^{I_{\ell+1}-2}\mathcal{P}_{a_{j+1}}(a_{|_{j}})\leq\sum_{i=a_{I_{\ell}}+1}^{\infty}X(a_{1}\cdots a_{I_{\ell}-1}i)\bigg\}=:\bigcap_{\ell=1}^{k}E_{a,\ell}.$
@@
[…]
@@
We now let $d_{j}:=I_{j+1}-I_{j}-1$ for $j\in[k-1]$ and $d_{k}:=m-I_{k}$ denote the number of entries between the running maxima of $a$. We can then define, for $(d_{j})_{j\in[k]}\in\mathbb{N}_{0}^{k}$ (with $0]:=\varnothing$),
$\mathscr{P}_{k}({}$ $a_{I_{1}},a_{I_{2}},\ldots,a_{I_{k}},d_{1},\ldots,d_{k})$ (3.15)
$=\{a\in\mathcal{U}_{\infty}:\text{ For all }j\in\{1,\ldots,k\}\text{ and all }i\in[d_{j}],\ a_{I_{j}+i}\in[a_{I_{j}}]\}$
@@
$\prod_{\ell=1}^{k}\!\!\!\!\sum_{\begin{subarray}{c}a^{(\ell)}:|a^{(\ell)}|=d_{\ell}+1\\ \text{$a^{(\ell)}$ $a_{I_{\ell}}$-conservative}\end{subarray}}\!\!\!\!\!\!\!\!\!\!\!\mathbb{P}\!\left(\widetilde{E}_{a,\ell}\right)\leq\prod_{\ell=1}^{k}\zeta^{d_{\ell}}\mathcal{M}_{\lambda_{a_{I_{\ell}}}}(Y_{a_{I_{\ell}}})\mathcal{L}_{\lambda_{a_{I_{\ell}}}}(\mathcal{P}_{a_{I_{\ell}}}).$ (3.18)
We can take out the factors $\zeta^{d_{\ell}}$ and use that the $d_{\ell}$ sum to mkm-k. Using this in (3.16) yields
$\sum_{m=1}^{\infty}{}$ $\sum_{k=1}^{m}\sum_{a_{I_{k}}>\ldots>a_{I_{1}}>K^{\prime}}\zeta^{m-k}\sum_{\begin{subarray}{c}(d_{j})_{j\in[k]}\in\mathbb{N}_{0}^{k}\\ \sum_{\ell=1}^{k}d_{\ell}=m-k\end{subarray}}\prod_{\ell=1}^{k}\mathcal{M}_{\lambda_{a_{I_{\ell}}}}(Y_{a_{I_{\ell}}})\mathcal{L}_{\lambda_{a_{I_{\ell}}}}(\mathcal{P}_{a_{I_{\ell}}})$ (3.19)
$\leq\sum_{m=1}^{\infty}\sum_{k=1}^{m}\binom{m-1}{k-1}\zeta^{m-k}\sum_{a_{I_{1}}>K^{\prime}}\cdots\sum_{a_{I_{k}}>K^{\prime}}\prod_{\ell=1}^{k}\mathcal{M}_{\lambda_{a_{I_{\ell}}}}(Y_{a_{I_{\ell}}})\mathcal{L}_{\lambda_{a_{I_{\ell}}}}(\mathcal{P}_{a_{I_{\ell}}})$
$=\sum_{m=1}^{\infty}\sum_{k=1}^{m}\binom{m-1}{k-1}\zeta^{m-k}\left(\sum_{n>K^{\prime}}\mathcal{M}_{\lambda_{n}}(Y_{n})\mathcal{L}_{\lambda_{n}}(\mathcal{P}_{n})\right)^{k}.$
@@
$\mathcal{U}_{L}:=\left\{u\in\mathcal{U}_{\infty}:u_{i}\leq L\text{ for all }i\in[|u|]\right\}\cup\{\varnothing\}.$
@@
Let $(\mathscr{T}_{t})_{t\geq 0}$ be a $(X,W)$-CMJ branching process satisfying Condition (c) of Assumption 2.1. Then, almost surely, for all $t\in(0,\infty)$, we have $|\mathscr{T}_{t}\cap\mathcal{U}_{L}|<\infty$ almost surely.
Recall that, for an $(X,W)$-Crump–Mode–Jagers branching process $(\mathscr{T}_{t})_{t\geq 0}$, we have
@@
Now, since $\tau_{\infty}<\infty$ almost surely, we infer from Proposition 3.3 with $L=K^{\prime}$, that $|\{u\in\mathcal{U}_{K^{\prime}}:\mathcal{B}(u)\leq\tau_{\infty}\}|<\infty$ almost surely. Therefore, by (3.21), the set
@@
The desired result follows by verifying the conditions on $\lambda_{n}$ and the vertex-weights in Corollary 2.7. We start by constructing a sequence $(\lambda_{n})_{n\in\mathbb{N}}$ such that there exists $N\in\mathbb{N}$ for which we have $\lambda_{n}<f(i,w)$ for all $i\geq n\geq N$. We recall that there exist $\varepsilon>0$, $n_{0}\in\mathbb{N}$, and a sequence $(k_{n})_{n\in\mathbb{N}}$ such that $\mathbb{P}\!\left(g(W)>k_{n}\right)\leq n^{-(1+\varepsilon)}$ for all $n\geq n_{0}$. Then, let $\delta\in(0,\varepsilon)$ as in the statement of Theorem 2.12, recall $\mu_{n}$ from (2.7), and set
@@
$\leq{}$ $\exp\bigg(\lambda_{n}\sum_{i=n}^{\infty}\frac{1}{s(i)-\lambda_{n}}-\frac{\lambda_{n}}{k_{n}}\sum_{i=0}^{n-1}\frac{1}{s(i)+\lambda_{n}/k_{n}}\bigg)+\exp\bigg(\lambda_{n}\sum_{i=n}^{\infty}\frac{1}{s(i)-\lambda_{n}}\bigg)\frac{1}{n^{1+\varepsilon}},$